# Theoretically Achieving Continuous Representation of Oriented Bounding Boxes

### v1

@@ 1 Introduction
Various models have been proposed for OOD [2, 18, 15, 7, 32]. However, as illustrated in Fig. 0(a) and Fig. 0(b), prevalent representation methods for OBBs exhibit discontinuity issues, encoding similar OBBs into distinct vectors. This introduces challenges in training neural networks as regression targets for similar input features may differ significantly, potentially causing confusion and hindering the training process.
The relationship between two OBBs can be conceptualized as one being transformed into the other through geometric operations: translation, rotation, scaling, and aspect ratio changes. While translation and scaling are relatively benign, rotation and aspect ratio changes are primary sources of discontinuity.
Rotation discontinuity, often referred to as the “Boundary Problem” [33] or “Rotation Sensitive Error” [21], stems from the periodicity of rotation angles. Although prior efforts [33, 35, 41] have addressed this, most of them still suffer from discontinuity arising from changes in aspect ratio. Some other methods, e.g. Gliding Vertex [32], aim to mitigate aspect ratio discontinuity but still struggle with rotation discontinuity. The discontinuity phenomena emerge during the encoding of OBBs into the regression target. Consequently, they can be termed encoding discontinuity.
@@ 2 Related Work
2.1 Oriented Object Detection (OOD)
With the increasing adoption of deep learning in computer vision, general object detection models [5, 9, 23, 13, 25, 14, 22] have emerged to enhance computers’ capacity for recognizing objects in natural images. Typically tailored for predicting HBBs, these models serve as the foundation for OOD when augmented with modules for OBB prediction. Rotated Faster R-CNN [23] stands as a prominent baseline for OOD, replacing HBB regression targets with OBBs. Several OOD methods, such as RoI Transformer [2], Gliding Vertex [32], ReDet [7], and Oriented R-CNN [30], follow a similar structure and can be implemented on the Rotated Faster R-CNN framework.
@@ 3 Theoretically Continuous Representation
Table 1: Comparison of Methods Dedicated to Discontinuity. Tar (R), Tar (A), Loss (R), and Loss (A) stand for Target Rotation Continuity, Target Aspect Ratio Continuity, Loss Rotation Continuity, and Loss Aspect Ratio Continuity, respectively. Dec (C) and Dec (R) stand for Decoding Decoding Completeness and Decoding Robustness. Acute-Angle is the common OBB representation that limits the rotation angle into a range of $\frac{\pi}{2}$ ($[-\frac{\pi}{4},\frac{\pi}{4})$ for example). Long-Edge is another common OBB representation, which determines the rotation angle $\theta$ by the long side and the x-axis, and the $\theta$ is within a range of length $\pi$.
@@ 3.1 Metrics for Continuity
To formally define continuity, we introduce $f_{Enc}$ as the mapping function from an OBB to $R^{n}$, and $f_{Dec}$ as the reverse mapping from a subset of $R^{n}$ to an OBB. Notably, $R(x,\theta)$ denotes the transformation generating an OBB $y$ by rotating the initial OBB $x$ by $\theta$ in a clockwise direction. Meanwhile, $A(x,r)$ generates a set of OBBs $\{y,z\}$ by adjusting one side of OBB $x$ to be $r$ times its original length. We refer to $S$ as the set of OBBs with areas larger than zero, and $L$ symbolizes the loss function. Models utilize $f_{Enc}$ to convert OBBs into regression targets and employ $f_{Dec}$ to translate prediction results into estimated OBBs.
@@
$\forall x\in S,\lim_{\theta\to 0}||f_{Enc}(x)-f_{Enc}\left(R(x,\theta)\right)||=0$ (1)
The corresponding discontinuity is often referred to as the ”Boundary Problem” [33, 35].
@@
$\forall x\in S,\lim_{\theta\to 0}||L\left(f_{Enc}(x)-f_{Enc}\left(R(x,\theta)\right)\right)||=0$ (3)
@@
$\forall x\in S,\lim_{r\to 1}\sum_{y\in A(x,r)}||L\left(f_{Enc}(x),f_{Enc}(y)\right)||=0$ (4)
@@
$\forall\epsilon>0,\forall x\in S,1-IoU(x,f_{Dec}(f_{Enc}(x)))<\epsilon$ (5)
@@
$\forall\epsilon>0,\exists\xi>0,\forall\Delta d\in R^{n}\land||\Delta d||<\xi,\forall x\in S,$ (6)
$1-IoU\left(x,f_{Dec}\left(f_{Enc}(x)+\Delta d\right)\right)<\epsilon$
@@ 3.2 Our Continuous Representation for OBB
$r_{s}=\begin{cases}\frac{x_{2}-x_{1}}{w}&w<h\\ \frac{y_{2}-y_{1}}{h}&w\geq h\\ \end{cases}$ (7)
@@
$y_{2}-y_{1}$ $=r_{s}h$ (8)
$x_{2}-x_{1}$ $=\frac{1-\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2}w$
@@ 3.3 Implementing COBB in OOD Models
$t_{x}$ $=\frac{\Delta x}{w_{p}},\ t_{y}=\frac{\Delta y}{h_{p}},\ t_{w}=\ln\left(\frac{w}{w_{p}}\right),\ t_{h}=\ln\left(\frac{h}{h_{p}}\right)$ (9)
@@
According to Eq. 7, the value of $r_{s}$ lies within the range [0, 0.5]. If an activation function, such as the Sigmoid function, is employed for $r_{s}$ prediction, the regression target for $r_{s}$ is computed as follows:
$r_{sig}=2r_{s}$ (10)
When no activation function is used to constrain the range of predicted $r_{s}$, the regression target is computed as follows:
$r_{ln}=\begin{cases}1+\log_{2}(r_{s})&r_{a}<0.5\\ 1+\log_{2}(1-r_{s})&r_{a}\geq 0.5\end{cases}$ (11)
@@
$s_{t}=(s_{0}^{\lambda},s_{1}^{\lambda},s_{2}^{\lambda},s_{3}^{\lambda})$ (12)
where $\lambda$ is set to 2 by default to amplify the gap between scores of the ground truth and other classifications.
@@ 3.5 Further Comparison with Peer Methods
Compared with Gliding Vertex: The Gliding Vertex method [32] represents an OBB by sliding the four vertices of its outer HBB. However, rotation continuity is compromised when the OBB is nearly horizontal, necessitating the utilization of an acreage ratio to prevent critical errors in such instances. Despite this precaution, Gliding Vertex still exhibits discontinuities in loss values and encoded results for horizontal cases. Moreover, its decoded results manifest as irregular quadrilaterals, and refining these into accurate OBBs introduces accuracy errors. In contrast, our methods ensure continuous prediction targets and loss values even for nearly horizontal OBBs, and the decoded quadrilaterals consistently represent accurate OBBs.
@@
Benchmark ArialDet OBBDet AlphaRotate MMRotate JDet
@@ 4.2 Components for Unified Benchmarking
  • Loss Function: Most models employ cross-entropy loss for classification and L1 loss for bounding box predictions. Some models use specially designed loss functions for OOD, such as IoU-Smooth L1 Loss [34].
@@ 4.3 Detection Models in the Benchmark
Experiments are conducted on multiple models within our benchmark framework, all subject to uniform conditions. The baseline models chosen for comparison were Rotated Faster R-CNN [23] and Rotated RetinaNet [13]. To minimize implementation discrepancies, other models are implemented with minimal alterations to their corresponding baseline architectures.
@@ 6 Conclusion
We have extensively shown the presence of boundary discontinuity in existing OOD models. We have presented a set of formal metrics, including Rotation Continuity, Aspect Ratio Continuity, Decoding Completeness, and Decoding Robustness, to quantify the continuity in existing methods. Employing these metrics, we have introduced COBB, an innovative continuous OBB representation method. Additionally, we standardized the implementation of mainstream models to establish an OOD benchmark using the JDet framework. Our experimental results on this benchmark showcase the effectiveness of our proposed method, achieving a notable improvement of 3.95% in $\text{mAP}_{75}$ on Faster R-CNN applied to the DOTA Dataset, without employing any additional techniques. COBB also has limitations. The outer HBB, sliding ratio $r_{s}$, and IoU scores exhibit irregular variations during OBB rotation, restricting its impact on rotation-equivariant detectors (e.g. ReDet [7]). Despite this, COBB proves effective in enhancing most OOD models by eliminating discontinuity.
@@ 7 Futher Explain of Metrics for Continuity
Within the Sec. 3.1, we introduced diverse metrics for evaluating methodological continuity. These metrics encompass Target Rotation Continuity, Target Aspect Ratio Continuity, Loss Rotation Continuity, Loss Aspect Ratio Continuity, Decoding Completeness, and Decoding Robustness. While formal definitions were provided in Sec. 3.1, this section delves deeper into their conceptual underpinnings.
@@
Loss Continuity: Loss Continuity encompasses two distinct components: Loss Rotation Continuity and Loss Aspect Ratio Continuity. This metric evaluates whether minor rotation or aspect ratio changes result in abrupt fluctuations in the loss value. While some methods might falter in maintaining target or aspect ratio continuity, they compensate by refining the loss function to ensure loss continuity. Notably, employing an L1 Loss function aids in preserving loss continuity for methods demonstrating target continuity.
Decoding Completeness: This criterion mandates precise representation of all OBBs. Methods rooted in CSL paradigms, such as those discussed in [33, 35, 27], often fall short in achieving Decoding Completeness due to discretized angle predictions, leading to imprecise orientation estimations of OBBs within finite angle classifications.
Decoding Robustness: Decoding Robustness demands that decoded OBBs remain resilient to slight errors in their representation. An example illustrating dissatisfaction with decoding robustness is GWD [36], which assigns square-like OBBs in different orientations to the same Gaussian distribution, leading to imprecise predictions for square-like targets. Notably, most OBB representation methods do not strictly satisfy decoding robustness, particularly in cases of minor disturbances to the rotation angle of OBBs with extremely large aspect ratios. Such corner cases are excluded from our analysis when evaluating OBB representation methods.
@@ 8.1 Derivation of Four OBBs with Identical Outer HBB and $r_{s}$
$\begin{cases}||P_{1}B||\cdot(h-||P_{1}B||)=h^{2}r_{s}(1-r_{s})\\ ||P_{1}A||\cdot(w-||P_{1}A||)=h^{2}r_{s}(1-r_{s})\\ \end{cases}$ (15)
@@
$\begin{cases}||P_{1}B||=\frac{1\pm(1-2r_{s})}{2}h\\ ||P_{1}A||=\frac{1\pm\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2}w\end{cases}$ (16)
@@ 8.2 OBB Recovery from Nine Parameters
$\begin{split}OBB_{0}=&[(x_{c}-x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}+y_{s}),\\ &(x_{c}+x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}-y_{s})]\\ OBB_{1}=&[(x_{c}+x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}+y_{s}),\\ &(x_{c}-x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}-y_{s})]\\ OBB_{2}=&[(x_{c}-x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}-y_{s}),\\ &(x_{c}+x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}+y_{s})]\\ OBB_{3}=&[(x_{c}+x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}-y_{s}),\\ &(x_{c}-x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}+y_{s})]\\ \end{split}$ (17)
@@
$y_{s}$ =$=\begin{cases}\frac{1-2r_{s}}{2}h&w\geq h\\ \frac{\sqrt{1-4\frac{w^{2}}{h^{2}}r_{s}(1-r_{s})}}{2}h&w<h\end{cases}$ (18)
$x_{s}$ =$=\begin{cases}\frac{\sqrt{1-4\frac{h^{2}}{w^{2}}r_{s}(1-r_{s})}}{2}w&w\geq h\\ \frac{1-2r_{s}}{2}w&w<h\end{cases}$

### v2

@@ 1 Introduction
Various models have been proposed for OOD [2, 20, 17, 7, 35]. However, as illustrated in Figs. 1(a)-1(b), prevalent representation methods for OBBs exhibit discontinuity issues, encoding similar OBBs into distinct vectors. This introduces challenges in training neural networks as regression targets for similar input features may differ significantly, potentially causing confusion and hindering the training process. The relationship between two OBBs can be conceptualized as one being transformed into the other through geometric operations: translation, rotation, scaling, and aspect ratio changes. While translation and scaling are relatively benign, rotation and aspect ratio changes are primary sources of discontinuity.
Rotation discontinuity, often referred to as the “Boundary Problem” [37] or “Rotation Sensitive Error” [23], stems from the periodicity of rotation angles. Although prior efforts [37, 39, 46] have addressed this, most of them still suffer from discontinuity arising from changes in aspect ratio. Some other methods, e.g. Gliding Vertex [35], effectively address aspect ratio discontinuity. Nonetheless, these techniques continue to face challenges in overcoming rotation discontinuity. The discontinuity phenomena emerge during the encoding of OBBs into the regression target. Consequently, they can be termed encoding discontinuity.
@@ 2 Related Work
2.1 Oriented Object Detection
With the increasing adoption of deep learning in computer vision, object detection models [16, 24, 30, 53, 5, 6, 11] have emerged to enhance computers’ capacity for recognizing objects in natural images. Typically tailored for predicting HBBs, these models serve as the foundation for OOD when augmented with modules for OBB prediction. Rotated Faster R-CNN [25] stands as a prominent baseline for OOD, replacing HBB regression targets with OBBs. Several OOD methods, such as RoI Transformer [2], Gliding Vertex [35], ReDet [7], and Oriented R-CNN [33], follow a similar structure and can be implemented on the Rotated Faster R-CNN framework.
@@ 3 Theoretically Continuous Representation
Table 1: Comparison of Methods Dedicated to Discontinuity. Tar (R), Tar (A), Loss (R), and Loss (A) stand for Target Rotation Continuity, Target Aspect Ratio Continuity, Loss Rotation Continuity, and Loss Aspect Ratio Continuity, respectively. Dec (C) and Dec (R) stand for Decoding Completeness and Decoding Robustness. Acute-Angle is the common OBB representation that limits the rotation angle into a range of $\frac{\pi}{2}$ ($[-\frac{\pi}{4},\frac{\pi}{4})$ for example). Long-Edge is another common OBB representation, which determines the rotation angle $\theta$ by the long side and the $x$-axis, and the $\theta$ is within a range of length $\pi$. Further explanation of this table is provided in the supplemental material, see Sec. 8.
@@ 3.1 Metrics for Continuity
To formally define continuity, we introduce $f_{Enc}$ as the mapping function from an OBB to $R^{n}$, and $f_{Dec}$ as the reverse mapping from a subset of $R^{n}$ to an OBB. Notably, $R(x,\theta)$ denotes the transformation generating an OBB $y$ by rotating the initial OBB $x$ by $\theta$ in a clockwise direction. Meanwhile, $A(x,r)$ generates a set of OBBs $\{y,z\}$ by adjusting one side of OBB $x$ to be $r$ times its original length. We refer to $S$ as the set of OBBs, and $L$ symbolizes the loss function. Models utilize $f_{Enc}$ to convert OBBs into regression targets and employ $f_{Dec}$ to translate prediction results into estimated OBBs.
@@
$\small\forall x\in S,\lim_{\theta\to 0}||f_{Enc}(x)-f_{Enc}\left(R(x,\theta)\right)||=0.$ (1)
The corresponding discontinuity is often referred to as the “Boundary Problem” [37, 39].
@@
$\small\forall x\in S,\lim_{\theta\to 0}||L\left(f_{Enc}(x),f_{Enc}\left(R(x,\theta)\right)\right)||=0.$ (3)
@@
$\small\forall x\in S,\lim_{r\to 1}\sum_{y\in A(x,r)}||L\left(f_{Enc}(x),f_{Enc}(y)\right)||=0.$ (4)
@@
$\small\forall x\in S,\exists t\in R^{n},IoU\left(x,f_{Dec}(t)\right)=1.$ (5)
@@
$\forall x\in S,\forall\epsilon>0,\exists\xi>0,\forall\Delta d\in R^{n}\land||\Delta d||<\xi,$ (6)
$1-IoU\left(x,f_{Dec}\left(f_{Enc}(x)+\Delta d\right)\right)<\epsilon.$
@@ 3.2 Our Continuous Representation for OBB
$\small r_{s}=\begin{cases}\frac{x_{2}-x_{1}}{w}&w<h,\\ \frac{y_{2}-y_{1}}{h}&w\geq h,\\ \end{cases}$ (7)
@@
$y_{2}-y_{1}$ $=r_{s}h,$ (8)
$x_{2}-x_{1}$ $=\frac{1-\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2}w.$
@@ 3.3 Implementing COBB in OOD Models
$t_{x}$ $=\frac{\Delta x}{w_{p}},\ t_{y}=\frac{\Delta y}{h_{p}},\ t_{w}=\ln\left(\frac{w}{w_{p}}\right),\ t_{h}=\ln\left(\frac{h}{h_{p}}\right),$ (9)
@@
According to Eq. 7, the value of $r_{s}$ lies within the range [0, 0.5]. To take advantage of this property, an effective way is to limit the range of prediction results, such as employing the sigmoid function. In this situation, the regression target for $r_{s}$ is computed as follows:
$\small r_{sig}=2r_{s}.$ (10)
Another method is extending the domain of $r_{s}$ as follows:
$\small r_{ln}=\begin{cases}1+\log_{2}(r_{s})&r_{a}<0.5,\\ 1+\log_{2}(1-r_{s})&r_{a}\geq 0.5.\end{cases}$ (11)
@@
$\small s_{t}=(s_{0}^{\lambda},s_{1}^{\lambda},s_{2}^{\lambda},s_{3}^{\lambda}),$ (12)
where $\lambda$ is a predefined constant to amplify the gap between scores of the ground truth and other classifications.
@@ 3.5 Further Comparison with Peer Methods
Compared with Gliding Vertex: The Gliding Vertex method [35] represents an OBB by sliding the four vertices of its outer HBB. However, rotation continuity is compromised when the OBB is nearly horizontal. Moreover, its decoded results manifest as irregular quadrilaterals, and refining these into accurate OBBs introduces accuracy errors. In contrast, our methods ensure continuous prediction targets and loss values for nearly horizontal OBBs, and the decoded quadrilaterals consistently represent accurate OBBs.
@@
Benchmark ArialDet OBBDet AlphaRotate [42] MMRotate [54] JDet
@@ 4.2 Components for Unified Benchmarking
  • Loss Function: Most models employ cross-entropy loss for classification and L1 loss for OBB prediction.
@@ 4.3 Detection Models in the Benchmark
Experiments are conducted on multiple models within our benchmark framework, all subject to uniform conditions. The baseline models chosen for comparison were Rotated Faster R-CNN [25] and Rotated RetinaNet [15]. To minimize implementation discrepancies, most of the other models are implemented with minimal alterations to their corresponding baseline architectures.
@@ 6 Conclusion
We have extensively shown the presence of boundary discontinuity in existing OOD models. To solve this problem, we have introduced COBB, an innovative continuous OBB representation method. Our experimental results showcase the effectiveness of our proposed method, achieving a notable improvement of 3.95% in $\text{mAP}_{75}$ on Rotated Faster R-CNN applied to the DOTA Dataset, without employing any additional techniques. COBB also has limitations. The outer HBB, sliding ratio $r_{s}$, and IoU scores exhibit irregular variations during OBB rotation, restricting its impact on rotation-equivariant detectors (e.g. ReDet [7]). Despite this, COBB proves effective in enhancing most OOD models by eliminating discontinuity.
7 Ackownledgement
The work was supported by the National Science and Technology Major Project under Grant 2021ZD0112902, the National Natural Science Foundation of China under Grant (62220106003, 62222607, 72342023), and the Research Grant of Beijing Higher Institution Engineering Research Center and Tsinghua-Tencent Joint Laboratory for Internet Innovation Technology.
@@ 8 Futher Explain of Metrics for Continuity
Within Sec. 3.1, we introduced diverse metrics for evaluating methodological continuity. These metrics encompass Target Rotation Continuity, Target Aspect Ratio Continuity, Loss Rotation Continuity, Loss Aspect Ratio Continuity, Decoding Completeness, and Decoding Robustness. While formal definitions were provided in Sec. 3.1, this section delves deeper into their conceptual underpinnings.
@@
Loss Continuity: Loss Continuity encompasses two distinct components: Loss Rotation Continuity and Loss Aspect Ratio Continuity. This metric evaluates whether minor rotation or aspect ratio changes result in abrupt fluctuations in the loss value. While some methods might falter in maintaining target rotation or aspect ratio continuity, they compensate by refining the loss function to ensure loss continuity. Notably, employing an L1 Loss function aids in preserving loss continuity for methods demonstrating target continuity.
Decoding Completeness: This criterion mandates precise representation of all OBBs. Methods rooted in CSL paradigms, such as those discussed in [37, 39, 29], often fall short in achieving Decoding Completeness due to discretized angle predictions, leading to imprecise orientation estimations of OBBs within finite angle classifications. Notably, we consider methods based on Gaussian distribution (such as GWD [40], KLD [41], and KFIoU [44]) satisfying Decoding Completeness because squares in various orientations can be possibly precisely decoded theoretically. However, in actual implementation, squares in only one orientation can be precisely decoded.
Decoding Robustness: Decoding Robustness demands that decoded OBBs remain resilient to slight errors in their representation. An example illustrating dissatisfaction with decoding robustness is GWD [40], which assigns square-like OBBs in different orientations to the same Gaussian distribution, leading to imprecise predictions for square-like targets. Especially, DHRec [22] encodes two symmetrical tilted slender OBBs into comparable representations, particularly when the aspect ratio is significantly large. Although this is uncommon, it suggests a propensity for the algorithm to confuse slender OBBs that are oriented in differing directions. Consequently, this observation leads us to conclude that DHRec does not fulfill the criteria for Decoding Robustness.
@@ 9.1 Derivation of Four OBBs with Identical Outer HBB and $r_{s}$
$\begin{cases}||P_{1}B||\cdot(h-||P_{1}B||)=h^{2}r_{s}(1-r_{s}),\\ ||P_{1}A||\cdot(w-||P_{1}A||)=h^{2}r_{s}(1-r_{s}).\\ \end{cases}$ (15)
@@
$\begin{cases}||P_{1}B||=\frac{1\pm(1-2r_{s})}{2}h,\\ ||P_{1}A||=\frac{1\pm\sqrt{1-4\cdot\frac{h^{2}}{w^{2}}\cdot r_{s}(1-r_{s})}}{2}w.\end{cases}$ (16)
@@ 9.2 OBB Recovery from Nine Parameters
$\begin{split}OBB_{0}=&[(x_{c}-x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}+y_{s}),\\ &(x_{c}+x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}-y_{s})],\\ OBB_{1}=&[(x_{c}+x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}+y_{s}),\\ &(x_{c}-x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}-y_{s})],\\ OBB_{2}=&[(x_{c}-x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}-y_{s}),\\ &(x_{c}+x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}+y_{s})],\\ OBB_{3}=&[(x_{c}+x_{s},y_{c}-0.5h),(x_{c}+0.5w,y_{c}-y_{s}),\\ &(x_{c}-x_{s},y_{c}+0.5h),(x_{c}-0.5w,y_{c}+y_{s})].\\ \end{split}$ (17)
@@
$y_{s}$ =$=\begin{cases}\frac{1-2r_{s}}{2}h&w\geq h,\\ \frac{\sqrt{1-4\frac{w^{2}}{h^{2}}r_{s}(1-r_{s})}}{2}h&w<h,\end{cases}$ (18)
$x_{s}$ =$=\begin{cases}\frac{\sqrt{1-4\frac{h^{2}}{w^{2}}r_{s}(1-r_{s})}}{2}w&w\geq h,\\ \frac{1-2r_{s}}{2}w&w<h.\end{cases}$